OpenStudy (anonymous):

Check my answer: Erika is flying a kite. If the length of string is 125 meters and the distance between where Erika is standing to the vertical distance from the kite is 75 meters, how high is the kite in meters? Is the answer, 100 meters?

OpenStudy (anonymous):

It's just pythagorus. a^2+b^2=c^2? It's going to be a right-angled traingle, with the string acting as the hypotenuse (c^2). The vertical distance is a or b. To work it out: 125^2=75^+b^2. Just work it out and square root the answer to find b.

OpenStudy (anonymous):